OpenStudy (hunter68):

A firecracker shoots up from a hill 145 feet high, with an initial speed of 80 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground?

OpenStudy (hunter68):

13 seconds 11 seconds Nine seconds Six seconds

OpenStudy (hunter68):

what I got so far is there...80 -16t^2+145t+

OpenStudy (hunter68):

@NetflixAndNo

OpenStudy (hunter68):

@rachethoes55

OpenStudy (hunter68):

@Anaise

OpenStudy (hunter68):

@ThatGuyAvery

OpenStudy (hunter68):

@ParthKohli

OpenStudy (hunter68):

@AloneS

OpenStudy (hunter68):

@peterg2001

OpenStudy (hunter68):

thanks @peterg2001

OpenStudy (peterg2001):

last one (D six seconds)

OpenStudy (hunter68):

i thought it was 6 just wasnt sure...

OpenStudy (anonymous):

v initial = 80 s initial = 145 H(t) = 0 (because we want to know when the height is back on the ground) 0 = −16t^2 + 80t + 145 quadratic formula t = 1/4 ( 10+- 7sqrt(5)) = -1.413 or 6.413 t = 6 seconds to get down x^2+3x-18=0 (x+6)(x-3)=0 x=-6,3

OpenStudy (anonymous):

Hope this helps